OpenStudy (anonymous):

Identify the x- and y-intercepts of f(x). Without graphing g(x), identify its x- and y-intercepts. f(x) = -x2 + 16 and g(x) = f(4x)

OpenStudy (anonymous):

x intercept=> y=0 => g(x)=0=f(4x)=-(4x)^2+16=0 y intercept = > x=0 => g(0)= f(0)= 16

OpenStudy (anonymous):

Okay! Awesome! Do you think you could help me with a couple more @divu.mkr

OpenStudy (anonymous):

yeah sure..

OpenStudy (anonymous):

Identify the x- and y-intercepts of f(x). Without graphing g(x), identify its x- and y-intercepts. f(x) = 4x + 12 and g(x) = (1/6)f(x)

OpenStudy (anonymous):

g(x)= (4x+12)/6 that even more simple..

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

then do it